OpenStudy (anonymous):

For question 4, find the volume of the square pyramid shown. Round to the nearest tenth if necessary. I'll draw a pic and show my work. Please help find what I'm doing wrong.

OpenStudy (anonymous):



OpenStudy (anonymous):

13 times 10=130

OpenStudy (anonymous):

So what I've done is taken the equation V= 1/3 B * h -simplified to- v=1/3 100 * 13 -simplified to- V= 1/3 (1300) -simplified to- 433.3. My answer choices are A)333.3 B)400 C)1200 D) 600

OpenStudy (anonymous):

But that's not right...